OpenStudy (anonymous):

Boyle's Law represents the relationship of volume as pressure is changed. What are the variables of Boyle's Law? pressure and number of molecules of gas volume and temperature pressure and temperature volume and number of molecules of gas pressure and volume

OpenStudy (anonymous):

pressure and volume

OpenStudy (ciarán95):

Boyle's Law states that: \[P \alpha \frac{ 1 }{ V }\] (i.e. The pressure is directly proportional to the inverse of the volume of the gas. This means that if we were to decrease the pressure, we would see a proportional increase in the volume of the gas, and vice versa.) Another way of writing this is to say that: \[P =k \frac{ 1 }{ V }\] or \[PV = k\] where k is some constant value.